OpenStudy (anonymous):

1. -9i - 17 = -26 2. 14.9 = 8.6 + 0.9m 3. 15w - 21 =-111

OpenStudy (anonymous):

Are you trying to solve for the variables ?

OpenStudy (anonymous):

yea i think

OpenStudy (anonymous):

I will do one for you just to show you.... 1. - 9i - 17 = - 26 First you want to get -9i by itself and everything else on the other side -9i - 17 = - 26 (add 17 to both sides) -9i - 17 + 17 = - 26 + 17 (combine like terms) -9i = - 9 (now you have to divide by -9 to get i by itself) -9i/-9 = -9/-9 i = 1 Do you understand how I did that ? Do you have any questions about what I just did ?

OpenStudy (anonymous):

Do you want to work on the 2nd one ?

OpenStudy (anonymous):

yeah so, 2.14.9 = 8.6 + 0.9m 8.6+8.6=0.9+8.6 0.9m=9.5 0.9m/0.9 0.9/0.9 m=1 ?

OpenStudy (anonymous):

2. 14.9 = 8.6 + .09m you want to get .09m on one side and everything else on the other side. Do you know what to do next ?

OpenStudy (anonymous):

my answer is above ur last response

OpenStudy (anonymous):

why are you adding 8.6 ?

OpenStudy (anonymous):

because the equation is 14.9=8.6+0.9m and 8.6 is as to 17 in the first equation

OpenStudy (anonymous):

To get .09m by itself, you have to get rid of the 8.6 on the right side. To do this you have to do the opposite of the + 8.6, which is subtract 8.6 from both sides. This will cause the 8.6 on the right side to cancel each other out. Does that make sense to you ?

OpenStudy (anonymous):

its 0.9 not .09 and yes

OpenStudy (anonymous):

Let me write it this way and see if it helps... 14.9 = 8.6 + .09m -8.6 -8.6 ------------ 6.3 = 0 + .09m does that make it easier to understand ?

OpenStudy (anonymous):

my bad change .09m to .9m

OpenStudy (anonymous):

so m=7?

OpenStudy (anonymous):

now we have : 6.3 = .9m (divide both sides by .9 to get m by itself) 6.3/.9 = .9m/.9 7 = m You got it correct. Way to go

OpenStudy (anonymous):

Shall we work on # 3 now ?

OpenStudy (anonymous):

yay ok let me try

OpenStudy (anonymous):

so w=-6?

OpenStudy (anonymous):

Let me check... 15w - 21 = - 111 + 21 + 21 ---------------- 15w + 0 = - 90 15w = - 90 15w/15 = - 90/15 w = - 6 BINGO.....yes it is correct. Very good :)

OpenStudy (anonymous):

yay i got it now

OpenStudy (anonymous):

By the way....I luv your user name....cool

OpenStudy (anonymous):

thnx nd i luv ur profile pic

OpenStudy (anonymous):

thanks.....I was glad to help you....I knew you could do it :)

OpenStudy (anonymous):

lol ok thnx

OpenStudy (anonymous):

i gave u a medal and became a fan

OpenStudy (anonymous):

awww thanks.....if you need any help, (if I can), just flag me down and I will try to help you.